Find the curvature of r(t) at the point (1, 1, 1). Given that r(t) = (t, t^2, t^3).
1
Expert's answer
2020-08-17T16:32:13-0400


Let's find the value for t where r(t) = P

t=1

curvate=|r'(t) x r"(t)|/|r'(t)|(x mean vector product of r'(t) and r"(t))

r'(t)=(1,2t, 3t2)

r''(t)=(0,2,6t)

r'(t) x r"(t)=(6t2 , -6t, 2)

|r'(t) x r"(t)|=(36t4+36t2+4)1/2

|r'(t)|=(1+4t2+9t4)1/2

|r'(t)|3=(1+4t2+9t4)3/2

curvate=(36t4+36t2+4)1/2/(1+4t2+9t4)3/2

Let's find curvate in t=1

curvate=(36+36+4)1/2/(1+4+9)3/2=761/2/143/2